OpenStudy (anonymous):

A board game uses a deck of 20 cards. Two cards are selected at random from this deck. Calculate the prob of selecting a monkey on the first draw and a frog on the second draw, both with replecement and without replacement?

OpenStudy (anonymous):

How many of each animal in the deck?

OpenStudy (anonymous):

5 monkeys, 5 frogs,5 birds and 5 lions

OpenStudy (dumbcow):

With replacement: 5/20 *5/20 = 1/16 Without replacement: 5/20*5/19 =5/76